OpenStudy (anonymous):

Solve the equation: 1/7(8t-9)=t-9/14

OpenStudy (anonymous):

www.wolframalpha.com

OpenStudy (anonymous):

enter your problem there. u will get steps along with answer

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

nvm

OpenStudy (the_fizicx99):

\[\frac{ 1 }{ 7(8t-9)} = \frac{ t-9 }{ 14 }\] I got \[t = \frac{ 9 }{ 2 }\]

OpenStudy (anonymous):

The answer is 9/5

OpenStudy (the_fizicx99):

Really..? 8t-9/7 = t -9/14 8t-9/7 = 14t/14 - 9/14 (common denominator to combine like terms) 8t-9/7 = 14t-9/14 Multiply them by a constant, 14(8t-9)/7 = 14(14t-9)/14 ... (14 divided by 7 = 2) 2(8t-9) = 14(14t-9)/14 (14's cancel out) 2(8t-9) = (14t -9) 16t - 18 = 14t - 9 -14t -14t 2t - 18 = -9 +18 +18 2t = 9 /2 /2 t = 9/2

OpenStudy (anonymous):

Alternate forms: 9/2 is correct. I just did equation on www.Wolframalpha.com